Please help
5/x+3=3/x-1

Answer by tutorcecilia(2152)

5/x+3=3/x-1 [Multiply both sides by the LCD=(x+3)(x-1)
(x+3)(x-1)5/x+3=(x+3)(x-1)3/x-1 [Cancel wherever possible]
(x-1)5=(x+3)3 [Multiply]
5x-5=3x+9 [Solve for x]
5x-3x=9+5
2x=14
x=7
.
Check by plugging (x=7) back into the original equation